Naruko885:

A watering can dispenses water at a rate of 0.40 gallon per minute. The original volume of water in the can was 10 gallons. Which set of ordered pairs shows the volume of water in the can in gallons (y) as a function of time in minutes (x) from the first minute after the can starts dispensing water? {(10, 1), (9.60, 2), (9.20, 3)} {(1, 10), (2, 9.60), (3, 9.20)} {(9.60, 1), (9.20, 2), (8.80, 3)} {(1, 9.60), (2, 9.20), (3, 8.80)}

Naruko885:

anyone have any ideas?

MARC:

The function should be like this \(f(x)=10-0.40x\)

MARC:

where x=time in minutes

MARC:

Keep in mind that,the value of x is the domain \(f(\color{red}{x})=10-0.40\color{red}{x}\)

MARC:

and f(x) is the codomain \(\color{blue}{f(x)}=10-0.40x\)

MARC:

when x=1 min, \(f(1)=10-0.40(1)\) \(f(1)=10-0.40\) \(f(1)=9.60\)

MARC:

when x=2 min, \(f(2)=10-0.40(2)\) \(f(2)=10-0.80\) \(f(2)=9.20\)

MARC:

when x=3 min, \(f(3)=10-0.40(3)\) \(f(3)=10-1.20\) \(f(3)=8.80\)

MARC:

ordered pair \(\{\color{red}{x},\color{blue}{y}\}\) where x is the domain and y is the codomain

MARC:

Hope this helps u... @Naruko885